# Diffusion Posterior Sampling for Nonlinear CT Reconstruction

### v1

@@ 3.2 Sparse-view CT Reconstruction
Figure 3 illustrates the flexibility of the DPS Nonlinear approach with application to a sparse-view CT reconstruction. Again, we see that the DPS Nonlinear approach yields an image with sharp edges and recognizable features. In this case, there is better agreement with the underlying ground truth (suggesting, perhaps, less reliance on the generative nature of DPS and a better conditioned reconstruction problem). The FBP reconstruction exhibits obvious streaking due to undersampling, and MBIR shows significant blur and noise. In terms of PSNR, our proposed method DPS Nonlinear yielded $59.54$ dB versus $27.66$ dB with FBP and $37.15$ dB with MBIR. For SSIM, our method achieved $0.6774$ while FBP and MBIR was $0.2392$ and $0.3093$, respectively. Again, these quantitative measures agree well with perceived image quality.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 3: Sparse-view CT reconstructions comparing, FBP, MBIR, and DPS Nonlinear. Colormap: W/L = 2000/0 HU.
@@
As in the low-dose situation, we generated 32 samples and computed a variance image. Three DPS Nonlinear reconstructions are shown to illustrate some of the variability associated with the stochastic reverse process. Three examples and the variance image are shown in Figure 4. Similar to the low-dose scenario, DPS Nonlinear tends to be more stable in the generation of low frequency general structures than the high frequency details in the lung and at tissue boundaries.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 4: Additional DPS sample reconstructions from the same data illustrating variability in the outputs.
@@ 3.3 OS-DPS Nonlinear Evaluations
Figure 5 illustrates the mean computation time, SSIM, and PSNR of 32 samples for OS-DPS Nonlinear with different numbers of ordered subsets. Note that the computation time reached a plateau when the OS = 24 at around 93 seconds with a peak at OS=1 at 106 seconds. This represents a speed-up of around $12\%$. Recall that the ordered subsets approach is only speeding up the construction of likelihood term in Equation (6), indicating that the reverse diffusion update in the DPS Nonlinear reverse process remain a significant contributor to overall computation time. We did not observe a significant drop on the SSIM and PSNR as the number of subsets increased. The mean SSIM went up $\sim 2\%$ and mean PSNR went down only $\sim 0.5\%$.
@@
Figure 5: The mean computation time, SSIM, and PSNR over the 32 samples generated under different number of order subsets with $1000$ iterations. The mean time cost represents the time OS-DPS Nonlinear takes to generate one sample.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 6: Comparison of sample mean image, bias image, and mean variance image under different order subsets. The first row of images are the sample mean image of different order subsets, the second row of images are the bias image under different order subsets, and the third row of images are the mean variance image of different order subsets. Colormap: W/L = 2000/0 HU for mean and bias image, W/L = 100000/50000 HU for standard mean variance image.
@@
To visualize OS-DPS reconstruction results, we chose to plot the mean, bias, and standard deviation images for a sampling of different numbers of ordered subsets (equal to 1, 6, 24, and 48). These results are shown in Figure 6. Again, the highest variability across all approaches lies in small lung features and at tissue boundaries. In general there appears to be an increase both in variability and bias as larger number of subsets are used. Interestingly, the six subset scenario appear that it might have some regional decreases in variability and bias relative to the single subset scenario, but this not evident in the previously computed PSNR (which is computed over the entire image).
Figure 7 demonstrates OS-DPS Nonlinear performance for only $200$ iterations for different numbers of ordered subsets. All of the results are generated in around 20 seconds, which is significantly faster than the $1000$ iteration case. Despite the limited computation time, the reconstructed data appear qualitatively similar to previous results. We found that OS-DPS Nonlinear achieved a PSNR of $42.48$ dB for 4 subsets, $43.83$ dB for 8 subsets, $44.33$ dB for 20 subsets. Similarly, the SSIM for 4 subsets was $0.3481$, 8 subsets was $0.3601$, and 20 subsets was $0.3730$. Though subtle in a subjective evaluation of the images, these quantitative results suggest that ordered-subsets can improve the image quality when computation time is limited.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Combined CT reconstructions comparing, 4 subsets, 8 subsets, and 20 subsets. Colormap: W/L = 2000/0 HU. These results are generated within only $200$ iterations of OS-DPS.
@@ 4 Discussion and Conclusion
We have demonstrated a new application of diffusion posterior sampling where a nonlinear measurement model is specified to combine prior knowledge of CT imagery from unsupervised training with a reverse process that is conditioned on the data likelihood model. We demonstrate the technique on two different measurement models to illustrate the application of DPS Nonlinear without the need for re-training for different measurement processes. We observe better performance than classic reconstruction methods; however, we also see the hallucinations that are possible with this and other generative approaches. How to address, analyze, and present data from generative methods remains an extremely important research area. In an investigation of the ordered-subsets variant of DPS Nonlinear, we found that OS-DPS Nonlinear does permit modest gains in terms of computational speedup and/or image quality (in a scenario of limited reconstruction time). We note that our implementation of DPS Nonlinear was careful to use projectors and backprojectors that are PyTorch compliant to reduce inefficiencies associated with GPU-CPU transfers or other data formatting issues. While these computational advantages may be enhanced with a more thorough optimization of all algorithm parameters and may be more significant for larger (e.g. 3D reconstruction) datasets; these results also suggest the need for acceleration of the computations associated with the reverse diffusion process.
In general, this initial study on DPS Nonlinear shows great promise; however, many aspects are not fully investigated. For example, a more comprehensive investigation of important algorithmic elements is required including selection of $\lambda$ and the required number of time steps, as these can have an impact on the balance between the prior and the data likelihood, as well as the output image quality.
While the preliminary data shown here illustrate the basic potential of the DPS approach, the prospects for DPS Nonlinear application across imaging systems is very broad. We continue to investigate applications where the nonlinear model is a critical element to reconstruction. For example, measurement models that exhibit significant detector blur are difficult (or impossible) to linearize - making the DPS Nonlinear approach an attractive option for reconstruction.

### v2

@@ 3.2 Sparse-view CT Reconstruction
Figure 5 illustrates the flexibility of the DPS Nonlinear approach with application to a sparse-view CT reconstruction. Again, we see that the DPS Nonlinear and DPS Linear yields an image with sharp edges and recognizable features. In this case, there is better agreement with the underlying ground truth (suggesting, perhaps, less reliance on the generative nature of DPS and a better conditioned reconstruction problem). The FBP reconstruction exhibits obvious streaking due to undersampling, and MBIR shows significant blur and noise. DOLCE (which was not trained for the sparse-view scenario) tends to have smoother features within the lungs and exhibits some significant false generative features including additional bones in the chest wall. In contrast, DPS Nonlinear (and Linear) appears to be much more robust in handling different measurement models. In terms of PSNR, DPS Nonlinear yielded $31.15$ dB versus DPS Linear with $31.29$ dB, and DOLCE with $24.18$ dB. For SSIM, DPS Nonlinear was $0.8451$ while DPS Linear and DOLCE was $0.8471$ and $0.7007$, respectively. These quantitative measures agree well with perceived image quality.
@@
Figure 5: Sparse-view CT reconstructions comparing FBP, MBIR (without and with TV regularization), DPS Linear, DOLCE, and the proposed DPS Nonlinear approach. All images share a common colormap with W/L = 2000/0 HU.
Figure 6 shows the root-mean-squared bias and standard deviation of DPS Nonlinear and DPS Linear for the sparse-view scenario. Again, there are similar performance curves, where both approaches achieve nearly the same bias and standard deviation optima. Balancing noise and bias, we chose an optimal value of $k=840$. We do observe a slight difference in mean bias of around $1.2\%$ for DPS Linear. However, we note that this improvement is within the ±$\pm~{}1.0$ standard deviation error bars (based on variability over DPS samples), suggesting that this difference is not particularly significant.
Figure 6: Bias and standard deviation comparison for sparse-view CT reconstructions comparing the proposed DPS Nonlinear and DPS Linear approach. Root-mean-squared bias and standard deviation are reported in Hounsfield units (HU) based on 16 generated samples from diffusion posterior sampling for each weighting coefficient $k$ value. Error bars (±$\pm~{}1.0$ standard deviation) are shown based on the variability over samples. The best weighting coefficient $k$ was selected to be $840$.
@@
As in the low-dose situation, we generated 16 samples and computed a variance image. Three DPS Nonlinear reconstructions are shown to illustrate some of the variability associated with the stochastic reverse process. Three examples and the variance image are shown in Figure 7. Similar to the low-dose scenario, DPS Nonlinear tends to be more stable in the generation of low frequency general structures than the high frequency details in the lung and at tissue boundaries.
@@
Figure 7: Additional DPS sample reconstructions from the same data illustrating variability in the outputs. Colormap: W/L = 2000/0 HU for sample image and W/L = 500/250 HU for standard deviation image.
3.3 Reconstruction across a range of fluence/noise levels
DPS Nonlinear CT reconstruction results under different fluence levels are shown in Figure 8. Among 400 DPS samples, we highlight the reconstruction that most closely achieves the mean PSNR in a local lung region of interest (ROI). As $I_{0}$ decreases, the mean ROI PSNR decreases from $27.57$ to $19.58$ dB. While the images appear sharp at the lowest exposures, one can see that the errors in the image are highly structured - appearing like typical lung features. Such hallucinations are more pronounced at the lowest exposure level. At the two highest exposures, errors are more consistent with the ground truth - appearing as differences in contrast (e.g. the feature is there but at a slightly different attenuation level). Moreover, the errors tend to be higher spatial frequency at higher fluence. Spatial variability is presented as a standard deviation map, where we see that the highest variability in samples occurs at features edges and in the texture of the lung, and is most pronounced at the lowest exposure.
Figure 8: DPS Nonlinear reconstructions from the measurement data with varying fluence levels from $I_{0}=100$ (noisiest) to $I_{0}=10^{5}$ (least noise). The top row shows reconstructed DPS samples that has closest PSNR to the mean PSNR of all $400$ generated samples. The second row shows the error map w.r.t the ground truth. The third row is the standard deviation map calculated based on $400$ samples. Colormap: W/L = 2000/0 HU for DPS Samples, W/L = 1000/0 HU for error map, W/L = 200/100 HU for std map.
The distributions of global and ROI PSNR results of $400$ samples under each fluence level are shown in Figure 9. The global and ROI PSNR trends are similar with better performance at higher noise levels. Because errors are more pronounced at edges and in high spatial frequency textures, the lung ROI PSNR values are lower than the global values. The distribution of performance across DPS samples is illustrated with ±$\pm~{}1.5$ standard deviation box, and the maximum and minimum performance is shown with the brackets/whiskers. Interestingly, the level of variation in PSNR is very similar across fluence levels. We see slightly more variability in performance across samples when looking at the lung ROI at the lowest fluence (e.g. the standard deviation is slightly larger). Similarly, the lowest $I_{0}$ shows some asymmetry in the local performance with a longer “tail” in the distribution for higher PSNR - suggesting a slight skew in the distribution toward higher performance.
To visualize maximum differences in performance, Figure 10 shows the error maps for the best and worst ROI PSNR for each fluence levels. The difference between best and worst is most pronounced for the lowest fluence level; however, significant errors are still present even in the best case. There are common errors across both fluence levels and within the sample distributions (notably for some horizontal features near the middle of the ROI). At higher fluence levels, errors appear to be more consistent between the best and worst case outputs.
Figure 9: Boxplot of global and ROI PSNR for $400$ samples generated under each fluence level. The orange middle line represents the mean PSNR. The box shows the ±$\pm~{}1.5$ standard deviation, and the top and bottom whiskers show the maximum and minimum PSNR.
Figure 10: Error maps for the best (top row) and worst (bottom row) case DPS samples in terms of ROI PSNR under different fluence levels. Colormap: W/L = 1000/0 HU for all images.
@@ 3.4 OS-DPS Nonlinear Evaluations
Figure 11 illustrates the mean cost time, SSIM, and PSNR of 16 samples for OS-DPS Nonlinear with different numbers of ordered subsets. Note that the computation time reached a plateau when the OS = 36 at around $333.09$ seconds with a peak at OS=1 at $371.05$ seconds. This represents a speed-up of around $10.24\%$. The general runtime of unconditioned DPS is also drawn on the plot with blue dotted line, which takes $330.28$ seconds. Recall that the ordered subsets approach is only speeding up the construction of likelihood term in Equation (6), indicating that the reverse diffusion update in the DPS Nonlinear reverse process remain a significant contributor to overall computation time. We did not observe a significant drop on the SSIM and PSNR as the number of subsets increased. The mean SSIM drop $\sim 5.6\%$ from $0.8017$ to $0.7563$ and mean PSNR drop $\sim 11.1\%$ from $27.56$ dB to $24.51$. Note that OS-DPS Nonlinear was demonstrated using LEAP version 0.8, while it cost around the same time (from $347.5$ seconds for OS=1 to $335.4$ for OS=180) for different ordered subsets under this low-dose scenario setting on LEAP version 1.0 and above. This speed-up was attribute to the CUDA optimization since LEAP version 1.0 where adaptive kernel was applied in single slice projection/backprojection, making the OS-DPS less sensitive to ordered subsets under this single projection setting. Nevertheless, the impact of ordered subsets on DPS will presumably be higher for larger data and volumes, especially fully 3D work in the future.
@@
Figure 11: The mean computation time, SSIM, and PSNR over the 16 samples generated under different number of order subsets with $1000$ iterations on DDPM sampler. Mean cost time is colored as blue dashed line with triangle markers, mean PSNR is colored as orange solid line with circle markers, mean SSIM is colored as green dot dashed line with square markers, time cost by the unconditioned DPS is colored as blue dotted line.
@@
Figure 12: Comparison of sample mean image, bias image, and mean variance image under different order subsets. The first row of images are the sample mean image of different order subsets, the second row of images are the bias image under different order subsets, and the third row of images are the mean variance image of different order subsets. Colormap: W/L = 2000/0 HU for mean and bias image, W/L = 1000/0 HU for bias image, and W/L = 500/250 HU for standard deviation image.
@@
To visualize OS-DPS reconstruction results, we chose to plot the mean, bias, and standard deviation images for a sampling of different numbers of ordered subsets (equal to 1, 6, 24, and 90). These results are shown in Figure 12. Again, the highest variability across all approaches lies in small lung features and at tissue boundaries. In general there appears to be an increase both in variability and bias as larger number of subsets are used.
Table 1 shows a runtime comparison using a single NVIDIA GeForce RTX 3090Ti GPU for image reconstruction. A comparison of various reconstruction approaches is presented, where we found that the proposed OS-DPS Nonlinear can perform a reconstruction in $333.09$ seconds, which is faster than DPS Nonlinear at $371.05$ seconds, DPS Linear at $370.12$ seconds, and DOLCE at $1199.80$ seconds. This illustrates that the ordered-subsets approach provides a modest improvement in runtime. Moreover, there does not appear to be a significant computation cost associated with the nonlinear model over the linearized approach. While we have not tried to optimize computation time for DOLCE, we find the DPS methods to be significantly faster based on our implementation of DOLCE.
Table 1: Runtime Comparison of single image reconstruction among different diffusion-based methods.
Methods Runtime (sec)
OS-DPS Nonlinear (OS=6) 333.09
DPS Nonlinear (OS=1) 371.05
DPS Linear 370.12
DOLCE 1199.80
@@ 4 Discussion and Conclusion
We have demonstrated a new application of diffusion posterior sampling where a nonlinear measurement model is specified to combine prior knowledge of CT imagery from unsupervised training with a reverse process that is conditioned on the data likelihood model. We demonstrated the technique on multiple systems to illustrate both the general applicability to different measurement models and noise levels, but also to show the advantages over current diffusion-based methods which rely on simplified (linearized) measurement models and/or conditional training of the diffusion network.
In an investigation of the ordered-subsets variant of DPS Nonlinear, we found that OS-DPS Nonlinear does permit modest gains in terms of computational speedup and/or image quality (in a scenario of limited reconstruction time). We note that our implementation of DPS Nonlinear was careful to use projectors and backprojectors that are PyTorch compliant to reduce inefficiencies associated with GPU-CPU transfers or other data formatting issues. While these computational advantages may be enhanced with a more thorough optimization of all algorithm parameters and may be more significant for larger datasets (e.g. 3D reconstruction where the projection operations are more costly); these results also suggest the need for acceleration of the computations associated with the reverse diffusion process.
Like other generative models, the proposed DPS Nonlinear technique is subject to hallucinations. We have illustrated such errors in an brief analysis of the distribution of DPS samples and found that the potential for these kinds of misrepresentation in the image increase in frequency and size as data fidelity/fluence decreases for DPS Nonlinear. While DPS grounds image reconstruction using a physics-based measurement model and provides a potential way to investigate hallucinations through multiple samples, users of such generative techniques should be aware of the potential for hallucination. How to address, analyze, and present data from generative methods remains an extremely important ongoing research area.
While the preliminary data shown here illustrates the potential of the DPS approach, the prospects for DPS Nonlinear application across imaging systems is very broad. We continue to investigate applications where the nonlinear model is a critical element to reconstruction. Examples include CT systems with projection blur from the detector[28], x-ray source, and/or gantry motion; as well as spectral CT[29] that combines reconstruction and material decomposition. Such scenarios can be difficult (or impossible) to linearize - making the DPS Nonlinear approach an attractive option for reconstruction.
Disclosures
The authors have no relevant financial interests or conflicts of interest to disclose.
Code, Data, and Materials Availability
Research code is currently not available. Research data is generated based on CT Lymph Nodes Dataset [21].
Acknowledgments
This work was supported, in part, by NIH grant R01CA249538. Thanks to Xiaoxuan Zhang (University of Pennsylvania, Philadelphia, PA, USA), Peiqing Teng (Johns Hopkins University, Baltimore, MD, USA).
Biographies and photographs of the other authors are not available.